# Accelerating Hamiltonian Monte Carlo for Bayesian Inference in Neural Networks and Neural Operators

### v1

@@ 4.1.2 Case II
In this example, the parameters for Eq. (21) used to generate the training and test data are: $\omega_{1}=4,\omega_{2}=-12,\phi_{1}=0,\phi_{2}=\pi/2,a=4,b=5,\sigma_{d}=0.05$. Again, twenty points with $x\in[-1,-0.2]\cup[0.2,1]$ are generated for training and 300 points with $x\in[-1.2,1.2]$ are generated for validation. In this case, a BNN with two hidden layers of 10 neurons each is considered with tanh activation functions, resulting in a total of 141 neural network parameters. The ELBO loss and MSE convergence during VI training are shown in Figure 5. Since the network is overparameterized, the training loss is much lower than the validation loss. Once the posterior distribution is obtained from VI, the sensitivities of the parameters are analyzed to determine the most sensitive parameters to quantify uncertainties in the outputs. The histogram of sensitivities for the parameters and the percentage of variance captured are shown in Figure 6. Here, the 79 parameters with the highest sensitivities are required to capture 90% of the total variance in the network prediction. Although the true function can be represented by a sine activation function with just six parameters, as seen in case I, the introduction of tanh as the activation function necessitates about 80 parameters to approximate the true function. This case demonstrates that a meaningful reduction in the parameter space is achieved even for a small network size, such that, in the VI-HMC step, the parameter space to sample from reduces to about 56% from the original space.
@@
Figure 6: Case II: BNN parameter sensitivities (a) Histogram of sensitivity values for the BNN parameters. (b) Percentage of captured variance for increasing number of BNN parameters.
@@
The predictions obtained by VI, HMC, and VI-HMC are compared in Figure 7. Compared to HMC, VI again overestimates the uncertainties where training data is present and underestimates the uncertainty in the interpolation region $x\in(-0.2,0.2)$ and the extrapolation regions $x\in[-1.2,-1)\cup(1,1.2]$. VI-HMC, on the other hand, provides uncertainties that better match the uncertainties from HMC. They are qualitatively better in their lower uncertainties near existing training data, while also yielding larger and more realistic uncertainties in the interpolation and extrapolation regimes. Further, since some parameters are fixed to be the means obtained from VI, the mean prediction of VI-HMC combines the mean predictions from VI and HMC. This effect can be seen in the interpolation regime $x\in(-0.2,0.2)$.
Figure 7: Case II: BNN predictions and their uncertainties from (a) VI, (b) HMC, and (c) VI-HMC.
samples, mean prediction, training data, true function.
@@ 4.2.1 Burgers’ equation
A periodic boundary condition is considered here such that
@@
The ELBO loss function and MSE during VI training are shown in Figure 8. A learning rate scheduler is used to reduce the learning rate to 10% of the current value when the validation loss plateaus. This results in the kinks observed near 6000 epochs in Figure 8. Following the VI training, parameter sensitivities are estimated as shown by the histogram in Figure 9(a). From Figure 9(b), we see that the number of parameters needed to capture 90% of the variance is only 28,214 out of the 172,401 total parameters. This corresponds to a substantial dimension reduction, resulting in significant computational savings. Without this reduction, HMC is computationally prohibitive, and therefore, results from full HMC are not shown here.
@@
Figure 9: Burgers’ Equation: Bayesian DeepONet sensitivities. (a) Histogram of sensitivity values for the Bayesian DeepONet (x-axis is limited to the 99th percentile of $S_{i}^{2}$ for clarity). (b) Ratio of variance captured for increasing number of Bayesian DeepONet parameters, $\dfrac{\sum_{i=1}^{n}S_{i}^{2}}{\sum_{j=1}^{N_{\Theta}}S_{j}^{2}}$ vs $n$
BNN predictions and the associated uncertainties are shown for the best and worst cases in Figure. 10. The mean predictions for the VI-HMC are reasonably accurate while showing well-correlated uncertainty estimates. The mean relative L2 errors, evaluated as $\frac{1}{N}\sum\frac{||Y-Y_{p}||_{2}}{||Y||_{2}}$, for VI and VI-HMC are 3.8% and 7.52% respectively.
Figure 10: Comparing predictions between VI (left) and VI-HMC (right) for the Burgers equation for 4 different initial conditions at time $t=1$. (a) Best prediction by VI, (b) Best prediction by VI-HMC, (c) Worst prediction by VI and (d) Worst prediction by VI-HMC samples, mean prediction, true solution.
@@ 4.2.2 Hypersonic flow over a cone
In the final example, a DeepONet is trained to learn the Navier-Stokes operator that maps the inlet conditions of a Mach 6, boundary layer flow over a cone to the wall pressure measurements. This problem is challenging due to the nonlinearity and chaotic nature of the Navier-Stokes operator. Specifically, in the regime of interest, the flow is in the laminar-to-turbulence transitional regime, and as such the downstream wall-pressure data are sensitive to small variational in the upstream disturbances [53, 54, 55, 56, 57]. Previous attempts to learn the Navier-Stokes operator in the hypersonic transitional regime have only considered deterministic networks [58, 59, 60]. In light of the extreme sensitivity to uncertainty, a Bayesian approach is desirable. The data is obtained from Morra et al [61]. It is generated by solving the compressible Navier-Stokes equation by direct numerical simulations. The input to the DeepOnet is the inlet condition described by its power spectra represented as $\mathbf{c}\in\mathbb{R}^{11}$. The output is the pressure measurement $\mathrm{p}$ at given sensor locations and the frequency of the pressure spectra.
The architecture of the DeepOnet is presented in Table. 3. The convergence of the ELBO loss for the VI training and the mean squared error loss during training is shown in Figure 11. Both losses show convergence after 300 epochs. The results of the sensitivity analysis are shown in Figure. 12. The total number of parameters in the DeepONet is 16,321. The sensitivity analysis shows that 840 parameters are enough to capture 90% variance. Thus, there is more than a 19 times reduction in the dimension of the parameter space to perform HMC, which accelerates HMC significantly.
@@
VI-HMC is performed to sample the parameters in the reduced space, and a comparison of uncertainties is shown in Figure. 13(h). As we saw for the simpler BNNs, VI overestimates uncertainties when compared to the hybrid VI-HMC approach.
@@
Figure 13: Flow over cone: Comparing uncertainties quantified by (a-d) VI and (e-h) VI-HMC.
training data DeepONet prediction, uncertainty (±$\pm 3\sigma$).
@@ 4.2.3 Computational cost
In the first approach, the integration of the Hamiltonian is performed using a fixed step size for a given problem when implementing the VI-HMC and the HMC methods. With the same step size, the acceptance rate, accuracy, and the time taken to draw one sample using the two approaches are compared in Table 2. The proposed VI-HMC method has a better acceptance rate while taking significantly less time compared to the HMC approach. Specifically for the Burgers equation, there is more than $2$x reduction in computational cost for the proposed VI-HMC approach compared to HMC, which can be attributed to the dimension reduction of the parameter space. In addition, the acceptance rate of the HMC approach is merely 10%, rendering the sampling inefficient in contrast to the 87% acceptance for VI-HMC. Similarly, for the cone problem, there is a $2.5$x reduction in computational cost for the VI-HMC method. Similar to the Burgers problem, the acceptance rate for HMC for the cone problem is just 55%, which implies poor sampling quality, compared to the 91% acceptance rate for the VI-HMC approach. Although theoretically, no time improvement per sample is expected as mentioned in Section 3.3, practical implementation of the full batch HMC results in gradient explosion, i.e, infinite values for the gradients of the parameters. Therefore, a symmetric splitting HMC is implemented following [25], which results in the increased computational cost for the two problems implemented in this work.
In the second approach, the acceptance rate for MCMC sampling is fixed at 80% and the step size is computed using the step size adaptation method proposed in Algorithm 5 of [62]. For the Burgers equation, the required step size for 80% acceptance rate with HMC is 5 times lower than the step size required for VI-HMC, which further demonstrates the computational advantage of the proposed approach. In the cone problem, the required acceptance step size for HMC is six orders of magnitude smaller than the step size required for VI-HMC, which renders the HMC approach infeasible to implement in an acceptable time frame.
Table 2: Comparison of the computation cost between HMC and VI-HMC for Bayesian Neural Operators. The step size is fixed at $42E-$ 4 and $51E-$ 5 for the Burgers and the cone problems, respectively (for column 3), and the acceptance rate is fixed at 80% for both problems (for column 4).
@@ 5 Limitations
Since the proposed method uses the VI technique in its initial training, many limitations in training a BNN with VI are carried over as limitations to the proposed method. This is especially noteworthy in cases where training VI and sampling from HMC require contradictory problem setups. For example, we found that for Case 2 of the BNN example, where the data is limited and the network is large, it is faster and more accurate to train a neural network with VI using a small standard deviation assumed for the Gaussian likelihood (noise in the data). In contrast, HMC struggles to converge to a posterior distribution that provides good predictions for a small standard deviation of Gaussian likelihood. On the other hand, when the noise standard deviation is high, HMC can converge faster while VI struggles to find the right distribution that predicts the given data. Since changing the likelihood or introducing posterior tempering alters the underlying distribution that is approximated or sampled, one should take utmost care in optimizing the hyperparameters such as the likelihood variance or the posterior temperature, when implementing the proposed VI-HMC method.
@@ 6 Conclusions
We propose a new hybrid approach that combines variational inference and Hamiltonian Monte Carlo for sampling from the posterior parameter distributions in Bayesian neural networks and neural operators. We demonstrate that different network parameters contribute to varying degrees in quantifying uncertainties in predictions and leverage this insight to formulate the hybrid approach. This approach uses sensitivity analysis on the posterior distributions learned from variational inference to significantly reduce the dimension of the parameter space. Hamiltonian Monte Carlo is then applied on this reduced parameter space, which greatly accelerates uncertainty quantification. The proposed approach provides accurate estimates of uncertainties, while being much faster in convergence than traditional HMC methods. We demonstrate the efficiency and accuracy of the framework by learning the posterior parameter distributions in both deep neural networks and operator networks and demonstrating their robust uncertainty estimates in the solution.
We demonstrated the application of the method to learning functions and operators. In the latter case, we demonstrated that we can accurately and efficiently predict the uncertainty of operator networks that represent the Burgers and Navier-Stokes equations. The present approach provides an important foundation for a number of future avenue, related to both forward and inverse problems. In the former context, the trained operator networks can be evaluated at very high speeds, and are therefore ideal for integration in control applications. In the latter context, specifically in data assimilation [63], the operator networks can accelerate the solution of the inverse problem by significantly reducing the cost of the forward model from that of solving the nonlinear differential equation to sample and evaluation of the network parameters.
@@ Appendix A DeepONet Architectures
Table 3: DeepOnet architecture for problems considered. Both trunk and branch networks are feed forward neural networks.
@@ Appendix B Sensitivity in the parameters
Parameters Sensitivities for the Bayesian DeepONets are visualized in Figure 14,15. In the Burgers example, the trunk net is less sensitive in predicting uncertainties than the branch net. The trunk net can be viewed as approximating the basis of the output function, and the branch net as approximating the coefficients of the output function. Parameters in the branch net being more sensitive to the outputs implies that uncertainties in the output are more sensitive to uncertainties in the coefficients than the basis functions. Further, in the trunk net of the burgers problem, the later layers (layers close to the outputs) are more sensitive than the initial layers (layers close to the inputs) which might be attributed to vanishing gradients in the initial layers.
In the cone problem however, the trunk network has more sensivitive parameters than the branch net indicating that the basis functions are more sensitive in this case. Further analysis is necessary to understand the spread of sensitivities across the network weights.
@@
Appendix C Hyperparameters of HMC and VI-HMC
For the HMC approach, the step size and the number of leapfrog steps are important hyperparameters that affect the sample quality. In all of the examples presented in this work, all hyperparameters except for the step size for HMC and VI-HMC approaches are set the same as shown in Table 4. The step size is manually tuned, and the number of steps is set to be $\frac{\pi\times\text{posterior variance}}{2\times\text{step size}}$. The posterior variance is taken as the maximum variance of parameters obtained from VI for the Burgers problem and the BNN’s case I and the median variance of parameters from VI for all other cases.
@@ Appendix C Hyperparameters of HMC and VI-HMC
Table 4: Hyperparameters for the HMC and VI-HMC approaches.
@@
Posterior variance $0.0679^{2}$ $0.2024^{2}$ $0.1027^{2}$ $0.0108^{2}$
Parallel chains [^3] A few bad chains are removed before post processing where applicable. 1 10 10 10

### v2

@@ 4.1.2 Case II
In this example, the parameters for Eq. (21) used to generate the training and test data are: $\omega_{1}=4,\omega_{2}=-12,\phi_{1}=0,\phi_{2}=\pi/2,a=4,b=5,\sigma_{d}=0.05$. Again, twenty points with $x\in[-1,-0.2]\cup[0.2,1]$ are generated for training and 300 points with $x\in[-1.2,1.2]$ are generated for validation. In this case, a BNN with two hidden layers of 10 neurons each is considered with tanh activation functions, resulting in a total of 141 neural network parameters. The ELBO loss and MSE convergence during VI training are shown in Figure 5. Since the network is over-parameterized, the training loss is much lower than the validation loss. Once the posterior distribution is obtained from VI, the sensitivities of the parameters are analyzed to determine the most sensitive parameters to quantify uncertainties in the outputs. The histogram of sensitivities for the parameters and the percentage of variance captured are shown in Figure 6. Here, the 79 parameters with the highest sensitivities are required to capture 90% of the total variance in the network prediction. Although the true function can be represented by a sine activation function with just six parameters, as seen in case I, the introduction of tanh as the activation function necessitates about 80 parameters to approximate the true function. This case demonstrates that a meaningful reduction in the parameter space is achieved even for a small network size, such that, in the VI-HMC step, the parameter space to sample from reduces to about 56% from the original space.
@@
Figure 6: Case II: BNN parameter sensitivities (a) Histogram of sensitivity values for the BNN parameters (x-axis is limited to the 99th percentile of $S_{i}^{2}$ for clarity). (b) Percentage of captured variance for an increasing number of BNN parameters.
@@
The predictions obtained by VI, HMC, and VI-HMC are compared in Figure 7. Compared to HMC, VI again overestimates the uncertainties where training data is present and underestimates the uncertainty in the interpolation region $x\in(-0.2,0.2)$ and the extrapolation regions $x\in[-1.2,-1)\cup(1,1.2]$. VI-HMC, on the other hand, provides uncertainties that better match the uncertainties from HMC. They are qualitatively better in their lower uncertainties near existing training data, while also yielding larger and more realistic uncertainties in the interpolation and extrapolation regimes. Further, since some parameters are fixed to be the means obtained from VI, the mean prediction of VI-HMC combines the mean predictions from VI and HMC. This effect can be seen in the interpolation regime $x\in(-0.2,0.2)$. In addition to these results, the robustness of the sensitivity scores for varying noise levels and training dataset size is discussed in D
Figure 7: Case II: BNN predictions and their uncertainties from (a) VI, (b) HMC, and (c) VI-HMC. samples, mean prediction, training data, true function.
@@ 4.2.1 Burgers’ equation
A periodic boundary condition is considered here, such that
@@
The ELBO loss function and MSE during VI training are shown in Figure 8. A learning rate scheduler is used to reduce the learning rate to 10% of the current value when the validation loss plateaus. This results in the kinks observed near 6000 epochs in Figure 8. Following the VI training, parameter sensitivities are estimated as shown by the histogram in Figure 9(a). From Figure 9(b), we see that the number of parameters needed to capture 90% of the variance is only 28,214 out of the 172,401 total parameters. This corresponds to a substantial dimension reduction, resulting in significant computational savings. Without this reduction, HMC is computationally prohibitive, and therefore, results from full HMC are not shown here. In addition, effect of varying the sensitivity threshold $\tau$ is shown in E
@@
Figure 9: Burgers’ Equation: Bayesian DeepONet sensitivities. (a) Histogram of sensitivity values for the Bayesian DeepONet (x-axis is limited to the 99th percentile of $S_{i}^{2}$ for clarity). (b) Ratio of variance captured for increasing number of Bayesian DeepONet parameters, $\dfrac{\sum_{i=1}^{n}S_{i}^{2}}{\sum_{j=1}^{N_{\Theta}}S_{j}^{2}}$ vs $n$ .
[…]
Figure 10: Comparing predictions between VI (left) and VI-HMC (right) for the Burgers equation for 4 different initial conditions at time $t=1$. (a) Best prediction by VI, (b) Best prediction by VI-HMC, (c) Worst prediction by VI, and (d) Worst prediction by VI-HMC samples, mean prediction, true solution, initial condition.
@@ 4.2.2 Hypersonic flow over a cone
In the final example, a DeepONet is trained to learn the Navier-Stokes operator that maps the inlet conditions of a Mach 6, boundary layer flow over a cone to the wall pressure measurements. This problem is challenging due to the nonlinearity and chaotic nature of the Navier-Stokes operator. Specifically, in the regime of interest, the flow is in the laminar-to-turbulence transitional regime, and as such, the downstream wall-pressure data are sensitive to small variations in the upstream disturbances [63, 64, 65, 66, 67]. Previous attempts to learn the Navier-Stokes operator in the hypersonic transitional regime have only considered deterministic networks [68, 69, 70]. In light of the extreme sensitivity to uncertainty, a Bayesian approach is desirable. The data is obtained from Morra et al [71]. It is generated by solving the compressible Navier-Stokes equation by direct numerical simulations. The input to the DeepOnet is the inlet condition described by its power spectra represented as $\mathbf{c}\in\mathbb{R}^{11}$. The output is the pressure measurement $\mathrm{p}$ at given sensor locations and the frequency of the pressure spectra.
The architecture of the DeepOnet is presented in Table 3. The convergence of the ELBO loss for the VI training and the mean squared error loss during training is shown in Figure 11. Both losses show convergence after 300 epochs. The results of the sensitivity analysis are shown in Figure 12. The total number of parameters in the DeepONet is 16,321. The sensitivity analysis shows that 840 parameters are enough to capture 90% variance. Thus, there is more than a 19 times reduction in the dimension of the parameter space to perform HMC, which accelerates HMC significantly.
@@
VI-HMC is performed to sample the parameters in the reduced space, and a comparison of uncertainties is shown in Figure 13. As we saw for the simpler BNNs, VI overestimates uncertainties when compared to the hybrid VI-HMC approach.
@@
Figure 13: Flow over cone: Comparing uncertainties quantified by (a-d) VI and (e-h) VI-HMC. training data DeepONet prediction, uncertainty ($\pm 3\sigma$).
@@ 4.2.3 Computational cost
In the first approach, the integration of the Hamiltonian is performed using a fixed step size for a given problem when implementing the VI-HMC and the HMC methods. With the same step size, the acceptance rate, accuracy, and the time taken to draw one sample using the two approaches are compared in Table 2. The proposed VI-HMC method has a better acceptance rate while taking significantly less time compared to the HMC approach. Specifically for the Burgers equation, there is more than 22x reduction in computational cost for the proposed VI-HMC approach compared to HMC, which can be attributed to the dimension reduction of the parameter space. In addition, the acceptance rate of the HMC approach is merely 10%, rendering the sampling inefficient in contrast to the 87% acceptance for VI-HMC. Similarly, for the cone problem, there is a 2.52.5x reduction in computational cost for the VI-HMC method. Similar to the Burgers problem, the acceptance rate for HMC for the cone problem is just 55%, which implies poor sampling quality, compared to the 91% acceptance rate for the VI-HMC approach. Although theoretically, no time improvement per sample is expected as mentioned in Section 3.3, practical implementation of the full batch HMC results in gradient explosion, i.e, infinite values for the gradients of the parameters. Therefore, a symmetric splitting HMC is implemented following [25], which results in the increased computational cost for the two problems implemented in this work.
In the second approach, the acceptance rate for MCMC sampling is fixed at 80% and the step size is computed using the step size adaptation method proposed in Algorithm 5 of [58]. For the Burgers equation, the required step size for an 80% acceptance rate with HMC is 5 times lower than the step size required for VI-HMC, which further demonstrates the computational advantage of the proposed approach. In the cone problem, the required acceptance step size for HMC is six orders of magnitude smaller than the step size required for VI-HMC, rendering the HMC approach infeasible to implement within an acceptable time frame.
The overall computational benefits of the proposed approach originate from two sources: reduced error accumulation in leapfrog steps and reduced gradient pathologies. The primary source is the reduced error accumulation in the leapfrog integration, resulting from the dimensionality reduction of the parameter space. This reduced error accumulation results in better acceptance rates when the step size is fixed and allows for larger step sizes when the acceptance rate is fixed, yielding computational savings. Additionally, the larger parameter size, combined with large step sizes in leapfrog integration, leads to gradient pathologies, which adds to the computational cost in HMC. Whereas our approach reduces such gradient pathologies, resulting in additional computational benefits in some cases. The breakdown of these benefits can be seen from Table 2. Improvements to the required step size at a fixed acceptance rate are a result of reduced error accumulation in the leapfrog integration. Whereas, a change in time/sample, if any, is a result of reduced gradient pathologies.
Table 2: Comparison of the computation cost between HMC and VI-HMC for Bayesian Neural Operators. The step size is fixed at 2E42E-4 and 1E51E-5 for the Burgers and the cone problems, respectively (for column 3), and the acceptance rate is fixed at 80% for both problems (for column 4).
@@ 5 Limitations
Since the proposed method uses the VI technique in its initial training, many of the limitations in training a BNN with VI are carried over to the proposed method. This is especially noteworthy in cases where training VI and sampling from HMC require contradictory problem setups. For example, we found that for Case 2 of the BNN example, where the data is limited and the network is large, it is faster and more accurate to train a neural network with VI by assuming a small standard deviation for the Gaussian likelihood (noise in the data). In contrast, HMC struggles to converge to a posterior distribution that yields good predictions under such a small likelihood standard deviation. On the other hand, when the likelihood standard deviation is large, HMC converges faster while VI struggles to identify a distribution that accurately predicts the given data. Since changing the likelihood or introducing posterior tempering alters the underlying distribution that is approximated or sampled, one must take utmost care in optimizing the hyper-parameters, such as the likelihood variance or the posterior temperature, when implementing the proposed VI-HMC method.
@@ 6 Conclusions
We propose a new hybrid approach that combines variational inference and Hamiltonian Monte Carlo for sampling from the posterior parameter distributions in Bayesian neural networks and neural operators. We demonstrate that different network parameters contribute to varying degrees in quantifying uncertainties in predictions and leverage this insight to formulate the hybrid approach. This approach uses sensitivity analysis on the posterior distributions learned from variational inference to significantly reduce the dimension of the parameter space. Hamiltonian Monte Carlo is then applied to this reduced parameter space, which greatly accelerates uncertainty quantification. The proposed approach provides accurate estimates of uncertainties while converging much faster than traditional HMC methods. We demonstrate the efficiency and accuracy of the framework by learning the posterior parameter distributions in both deep neural networks and operator networks and demonstrating their robust uncertainty estimates in the solution.
We demonstrated the application of the method to learning functions and operators. In the latter case, we demonstrated that we can accurately and efficiently predict the uncertainty of operator networks that represent the Burgers and Navier-Stokes equations. The present approach provides an important foundation for a number of future avenues, related to both forward and inverse problems. In the former context, the trained operator networks can be evaluated at very high speeds and are therefore ideal for integration in control applications. In the latter context, specifically in data assimilation [72], the operator networks can accelerate the solution of the inverse problem by significantly reducing the cost of the forward model. This higher cost in the forward model stems from solving the nonlinear differential equation, and this cost is reduced by inexpensive sampling of the network parameters and evaluating the outputs.
@@ Appendix A DeepONet Architectures
Table 3: DeepOnet architecture for problems considered. Both trunk and branch networks are feed-forward neural networks.
@@ Appendix B Sensitivity in the parameters
Parameters Sensitivities for the Bayesian DeepONets are visualized in Figure 14,15. In the Burgers example, the trunk net is less sensitive in predicting uncertainties than the branch net. The trunk net can be viewed as approximating the basis of the output function, and the branch net as approximating the coefficients of the output function. Parameters in the branch net being more sensitive to the outputs implies that uncertainties in the output are more sensitive to uncertainties in the coefficients than the basis functions. Further, in the trunk net of the burgers problem, the later layers (layers close to the outputs) are more sensitive than the initial layers (layers close to the inputs), which might be attributed to vanishing gradients in the initial layers.
In the cone problem, however, the trunk network has more sensitive parameters than the branch net, indicating that the basis functions are more sensitive in this case. Further analysis is necessary to understand the spread of sensitivities across the network weights.
@@
Appendix C Hyper-parameters of HMC and VI-HMC
For the HMC approach, the step size and the number of leapfrog steps are important hyper-parameters that affect the sample quality. In all of the examples presented in this work, all hyper-parameters except for the step size for HMC and VI-HMC approaches are set the same values as shown in Table 4. The step size is manually tuned, and the number of steps is set to be $\frac{\pi\times\text{posterior variance}}{2\times\text{step size}}$. The posterior variance is taken as the maximum variance of parameters obtained from VI for the Burgers problem and the BNN’s case I, and the median variance of parameters from VI for all other cases.
@@ Appendix C Hyper-parameters of HMC and VI-HMC
Table 4: Hyper-parameters for the HMC and VI-HMC approaches.
@@
Posterior variance $0.0679^{2}$ $0.2024^{2}$ $1027^{2} 0.0$$0108^{2}$
Parallel chains [^3] A few bad chains are removed before post-processing, where applicable. 1 10 10 10
@@
Appendix D Robustness of sensitivity scores
The sensitivity scores are robust when the mean and standard deviation of the BNN parameters are unchanged for different training scenarios. When the choice of the dataset or the noise level affects the distribution of parameters learned by the network, the sensitivity scores are altered. This is expected since the sensitivity scores (in Eq.17) rely on the gradient of predictions with respect to parameters and the standard deviation of parameters. Any change in these gradients or standard deviation results in different sensitivities of the parameters, resulting in different contributions from the individual parameters to the total sensitivities. To verify this, experiments with different datasets and noise levels are performed for the neural network case II, and the results of the sensitivity analysis are presented.
Varying the noise level changes the mean and significantly changes the standard deviation of parameters learned by the Bayesian neural network, whereas varying the number of training data points doesn’t alter the distribution of parameters as much, as seen in Figure 16. As a result of this, the ratio of variance captured is more sensitive to varying noise level than varying the training data, as seen in Figure 17.
(a) Varying noise level
(b) Varying training data
Figure 16: Kernel Density estimates of the mean and standard deviation of BNN parameters learned for varying noise level $\sigma$ and number of training data points $N_{d}$. $\sigma=0.1,N_{d}=10$; $\sigma=0.05,N_{d}=20$; $\sigma=0.01,N_{d}=30$; $\sigma=0.005,N_{d}=40$
(a) Varying noise level
(b) Varying training data
Figure 17: Ratio of variance captured for different noise level $\sigma$ and number of training data points $N_{d}$. $\sigma=0.1,N_{d}=10$; $\sigma=0.05,N_{d}=20$; $\sigma=0.01,N_{d}=30$; $\sigma=0.005,N_{d}=40$
Appendix E Effect of the sensitivity threshold
We conducted experiments with varying sensitivity thresholds, $\tau$, to analyze the performance of the proposed approach on the Burgers dataset. With decreasing $\tau$, the number of parameters considered to perform HMC, i.e., the dimensionality of the parameter space, reduces as expected. Due to this dimensionality reduction, the computational cost to perform HMC also reduces, as summarized in Table 5. This is evident from the acceptance rate for a fixed step size (column 3) and the step size required to perform HMC for a fixed acceptance rate (column 4). The accuracy slightly improves since more parameters are fixed at mean values of VI with reducing $\tau$. Since the least sensitive parameters are fixed at their mean values, the uncertainties in the prediction also reduce with decreasing $\tau$ as seen in Figure 18. The uncertainties are also compared with randomly choosing 1,279 parameters (the same number of parameters as $\tau=0.3$) to elucidate the importance of the proposed sensitivity analysis. Importantly, it is evident that the uncertainties in the case of randomly choosing a subset of parameters are much worse compared to the uncertainties obtained by the parameters chosen from the sensitivity analysis, even with a small threshold of 0.3.
Table 5: Comparison of the computation cost between varying sensitivity threshold $\tau$ for the Burgers problem. The step size is fixed at 2E42E-4 (for column 3), and the acceptance rate is fixed at 80% (for column 4).
$\mathbf{\tau}$ # of parameters Fixed step size Fixed acceptance rate
Acceptance rate MSE Time/sample Required step size
0.9 28,214 87% 2.9E-4 6.27 s 5.0E-4
0.75 12,944 89% 2.9E-4 6.37 s 6.9E-4
0.50 4,007 92 % 2.7E-4 6.21 s 7.7E-4
0.30 1,279 94 % 2.0E-4 6.05 s 8.9E-4
(a) $\tau=0.9$
(b) $\tau=0.3$
(c) Random
(d) Uncertainty
Figure 18: Comparing predictions for two values of the sensitivity threshold and a randomly chosen subset of the parameters: (a-c) samples, mean prediction, true solution. (d) Comparing uncertainties for the three cases $\tau=0.9$ $\tau=0.3$, Random